Using one of the laws of exponents, prove that any number raised to the power 0 is 1.
Answer by stanbon(48568)

Let "x" be any number except zero.
Then x/x = 1
But
x/x = x^1/x^1 = x^(1-1) = x^0
Therefore x^0 = 1 for all numbers except x=0
=================
Cheers,
Stan H.